### v1

@@ 3.1 Fairness Terminology
In this subsection, we introduce and define some of the concepts and terminologies generally used in algorithmic fairness research that are relevant to this paper.
@@
  • Disparate impact ratio (DIR): This is a fairness metric which compares the rate of positive outcomes in the unprivileged class, by the rate of positive outcomes in the privileged class, as shown in equation 1. It is the measure of how different outcomes are for different groups, based on the results of a model Feldman et al. (2015). A value of less than 1 indicates less favorable outcomes for the unprivileged group. In this paper, we assume an acceptable lower bound of 0.8, and an upper bound of 1.2, with 1 being the ideal score.
@@
  • Equalized odds: This fairness metric enforces that the model correctly identify the positive outcome at equal rates across groups, and misclasify the positive outcome at equal rates across groups (creating the same proportion of True Positives and False Positives across groups). In this work, we represent this metric using 2 values; Difference in False Negative scores, and Difference in False Positive scores.
3.2 Multi-task Learning Based Bias Mitigation Method (Proposed Method)
Our proposed methodology is based on the premise of epistemic uncertainty in Bayesian uncertainty estimation. Uncertainty estimation refers to estimating the confidence level of a model output, and epistemic uncertainty refers to uncertainty in the structure and parameters of a model caused by a lack of knowledge, which has been proven to correlate with model weights Li et al. (2021); Wang et al. (2021b). According to the Bayesian approach, the posterior distribution of model weights can theoretically be obtained and used to make inferences by fitting a defined model to the observed data (training set). In this case, the posterior variance represents the total uncertainty of the prediction Wang et al. (2021b).
We hypothesize that when the model is most uncertain about the protected label, the weights of that model will lack knowledge at certain regions of the feature space related to the protected label. This means that using these weights for target label prediction will minimize the influence of that protected label on the final prediction. Based on this premise, our proposed method will minimize any bias introduced to the model regarding the protected label.
Figure 1 represents a visual diagram of our proposed method, which we described in our initial work (Zanna et al. (2022)). We utilize multi-task learning to predict the final prediction label and one of the protected labels (e.g., gender). We save the model periodically in our experiments and implement Monte Carlo dropout, allowing us to get uncertainty estimations of these saved models.
@@ 3.2 Multi-task Learning Based Bias Mitigation Method (Proposed Method)
3.3 Monte-Carlo Dropout
Modeling uncertainty with Monte Carlo dropout works by running multiple forward passes through the model with a different dropout mask every time. It draws weights from approximate posterior distributions. The posterior distribution is a way to summarize what we know about uncertain quantities in Bayesian analysis. It combines the prior distribution and the likelihood function, which tells you what information is in your observed data (the “new evidence”). In other words, the posterior distribution summarizes what we know after the data is observed.
@@ 3.3 Monte-Carlo Dropout
More information on this can be found in the original publication on Monte Carlo dropout by Gal et. al Gal and Ghahramani (2016).
@@ 3.4 Non-Dominated Sorting
From the distribution of Monte Carlo models obtained, we saved their uncertainty scores and extracted the parameters of each model to utilize in a single (non-multi-task) prediction of our main target label. We trained each of these new models, collected their performance scores, and calculated the fairness scores (DIRs, difference in false negative scores, difference in false positive scores) from their produced predictions.
We implemented a non-dominated sorting algorithm for the performance scores and uncertainty scores, specifically the DIR scores of the model ensemble, to obtain the Pareto optimal set of models regarded as the "Pareto Front", that give the ideal trade-offs of performance and fairness. With the Pareto Front (or frontier), there are no better performing models in terms of performance and fairness scores, and there are only models where an improvement in one score leads to a reduction in the other. Non-dominated sorting allows one to search for optimal solutions for an optimization problem with multiple objectives. The algorithm in Figure 2 describes this process of non-dominated sorting. This algorithm takes two equally-sized lists, one is a list of performance scores for the models produced by the implementation of Monte Carlo dropout represented by X, and the second is fairness scores, specifically the DIR scores of said models, represented by Y. This method returns the elements which lie on the Pareto frontier, sorted in order. In our case, the objective is to maximize X, which is the performance score, and minimize the absolute difference between Y (DIRs), and 1.
Figure 2: Algorithm to Calculate Pareto Front
4 Data
@@
We evaluated the effectiveness of our proposed method using three datasets, ADULT (individual income and demographic information) Kohavi et al. (1996), MIMIC-III (data from patients in critical care units) Johnson et al. (2016), and SNAPSHOT datasets (multi-modal physiological, behavioral, and survey data collected from 350 college students) Sano et al. (2015). See the Appendix A for more the details on the datasets.
4.1 ADULT Dataset
This dataset contains 14 categorical and continuous integer attributes from which we dropped columns with noisy data (over 100 missing values). We were left with 8 features: Age, Workclass, Education, Marital-status, Occupation, Relationship, Race, and Sex. Appendix A describes these features and their classes.
The prediction task for this dataset is a binary income/salary prediction, where a label of 1 represents participants who earn $>$ 50k and 0 for those who earn <=<=< = 50k in a year. We converted all the string features to integers, and we used 45222 of 48842 instances after dropping a total of 3,620 samples with noisy data. We utilized age, sex (gender), and race as the protected labels in our experiments.
For fairness analysis, we binarized the 3 protected labels to 0 as unprivileged and 1 as privileged classes. For the age label, 12% of the female class earned more than 50k in salary, while for the male class, 32% earned more than 50k. 36% of participants above the age of 40 earned more than 50k, while 17% of those below the age of 40 earned more than 50k. As for the race label, less than 13% of the participants who identified as Black, Amer-Indian-Eskimo, and "Other" earned more than 50k, while about 30% of those who identified as White or Asian-Pacific -Islander earned more than 50k. We chose the class corresponding to higher income/salary as the privileged class for each protected label. These privileged classes are also the majority classes in the dataset. For the age label, since the difference in the number of participants between classes is larger than the difference in average salary earned, we chose the class with more participants (less than or equal to 40) as the privileged class (1), while the class with fewer participants (over 40) as the unprivileged class (0). For race, Black(0), Amer-Indian-Eskimo(4), and other(2) became 0, and White(3) and Asian-Pac-Islander(1) became 1, and for sex, male became 1 and female became 0 as seen in table 1. We excluded the protected labels as features to feed into our prediction model. We split the data into 75% training and 25% testing sets to fit into the model.
@@ 4.1 ADULT Dataset
Table 1: ADULT: Binarization of Protected labels
Label Privileged Class (1) Unprivileged Class (0)
Age Less than or equal to 40 years Greater than 40 years
@@
4.2 MIMIC-III Dataset
With this dataset, we predicted in-hospital mortality (IHM) as a binary label. We determined IHM by comparing the patient’s date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays). We processed data by reproducing the method in Harutyunyan et al. (2019). A full description can be found in Appendix A and the original publication Harutyunyan et al. (2019). To process the data for fairness evaluation, we binary-encoded each protected label and assigned them as privileged and unprivileged classes to analyze them using our chosen fairness metrics. As shown in Table 2, we encoded all participants with age $<=60$ as the unprivileged class (0), and those $>60$ as the privileged class (1). The reasoning behind this encoding is that the Centers for Disease Control and Prevention (CDC) defines an “older adult” as someone who is at least 60 years old for Disease Control et al. (2012). For the Gender label, male was encoded as the privileged class (1) due to its higher number of samples and because female patients appear to have a slightly higher IHM risk, making the female class the unprivileged class (0). We classified the participants according to whether they have private or government-assisted insurance. Those with private or self-care insurance represented the privileged class (1), and those with Medicare, Medicaid, and Government represented the unprivileged class (0). The reasoning behind this is the connection between government-aided insurance and socioeconomic conditions and age as mentioned earlier in this section. For marital status, we categorized married patients, those with life partners as privileged, and all others (no partner) as unprivileged. We did this in accordance with research showing that unmarried patients, including those who are divorced, separated, widowed, or never married, show elevated rates of certain health complications and death compared to those with partners Dhindsa et al. (2020); Johnson et al. (2000).
Table 2: MIMIC-III: Binarization of Protected labels
Label Privileged Class (1) Unprivileged Class (0)
Age Greater than 60 years Less than or equal to 60 years
@@ 4.2 MIMIC-III Dataset
4.3 SNAPSHOT Dataset
[…]
Some participants do not have race and ethnicity information, and we dropped samples for such participants when running experiments for race and ethnicity labels. In the end, we dropped 314 samples from the training data, reducing it from 4030 to 3716 samples, and the test samples were reduced from 1866 to 1715. We split the data into 70% training and 30% testing sets for all experiments to fit into the model.
Figure 3: SNAPSHOT Demographic Distribution
5 Experiments
@@ 5.1 Baseline Models and Fairness Analyses
The first part of our experiments involved developing baseline models that make the target predictions for all datasets (income for ADULT, IHM for MIMIC-III, and 4 happiness and calmness labels for SNAPSHOT). Next, we analyzed our prediction results for fairness by calculating the DIR and equalized odds for predictions against the protected labels for all datasets, respectively. We conducted these analyses to determine whether and where the models were introducing bias and ensure that we only apply our bias mitigation method to aspects of the data that are actually biased.
For the ADULT dataset, we built a Convolutional Neural Network (CNN) with 2 dense layers and a dropout later with a dropout rate of 0.25 to predict binary income labels; $<=50$k as 0 and $>50$k as 1. We used a 3-fold cross-validation method and binary cross entropy as the loss function with an Adam optimizer. The structure of this model is shown in Appendix B
For the MIMIC-III dataset, we used a Long-Short Term Memory (LSTM)-based model proposed in Harutyunyan et al. (2019). LSTM is a Recurrent Neural Network (RNN) designed to capture long-term dependencies in sequential data Harutyunyan et al. (2019). An RNN is a type of Artificial Neural Network (ANN) with recurrent connections capable of modeling sequential data for sequence recognition and prediction Bengio et al. (1994). To implement our model, we resampled the time series data into regularly spaced intervals. In the case of multiple measurements of the same variable in the same interval, we utilized the value of the last measurement. We used a 48-hour window for prediction, enabling the detection of patterns that may indicate changes in patient acuity, as proposed by Harutyunyan et al. (2019). We selected the channel-wise LSTM model without deep supervision for our study based on its superior reported area under the receiver operating characteristic (AUROC) performance among the five developed non-multitask models by Harutyunyan et al. Harutyunyan et al. (2019). AUROC is a performance metric that demonstrates a model’s ability to discriminate between cases (positive examples) and non-cases (negative examples.) Hanley and McNeil (1982). We also utilized the area under Precision-Recall (PR) curve (AUPRC), a single number that summarizes the information in the PR curve. AUPRC is a useful performance metric for imbalanced data in a problem setting where finding positive examples is essential. Particularly in medical datasets, the AUPRC is frequently smaller in absolute value than the AUROC since the fraction of positives is often less than 0.5, meaning that AUPRC has a lower baseline value than AUROC Ozenne et al. (2015).
This selected model is a modified version of the standard LSTM in which all variables are first independently pre-processed with an individual bidirectional LSTM layer instead of working directly on the entire matrix of clinical events as usual. Bidirectional LSTM (BiLSTM) is a recurrent neural network where, unlike standard LSTM, the input flows in both directions and can utilize information from both sides Graves and Schmidhuber (2005). Our network consists of a bi-directional layer, an LSTM layer with 16 neurons, and a dropout layer with a dropout rate of 0.3 (structure shown in Appendix B). We used a batch size of 8, binary cross entropy as the loss function and Adam optimizer. We executed this for 100 epochs, reproducing the experiments from Harutyunyan et al. (2019).
The baseline model we used for predicting the labels for the SNAPSHOT dataset is a GRU model with 2 GRU layers and 2 dropout layers with a dropout rate of 0.5, 30 neurons and focal loss with a gamma value of 2, and an alpha value of 4. This model predicts 4 labels, namely: "morning happiness", "morning calmness", "evening happiness", and "evening calmness". We trained this model for 300 epochs with a batch size of 1000 and a focal loss function since there is a significant class imbalance in the target label for this dataset. Focal loss is an extension of the cross-entropy loss function that would down-weight easy examples and focus training on hard negatives Lin et al. (2017). So it reduces the loss contribution from easy examples and increases the importance of correcting misclassified examples. The structure of this model is also shown in Appendix B. With 8 protected labels, which are gender, race, ethnicity, openness, conscientiousness, extraversion, agreeableness, and neuroticism, and 4 target labels, we conducted 32 sets of experiments to analyze the bias in this dataset.
After performing fairness analysis on each of baseline model, we selected the protected labels by which our models showed the most bias and tested our methods on all datasets. For all fairness analyses in our experiments, we utilized a fairness toolkit developed by IBM - AI Fairness 360 Bellamy et al. (2019).
@@ 5.2 Proposed Method Implementation
After running the multi-task networks and obtaining an ensemble of models using the parameters described in Appendix B, we calculated the uncertainties of these models. We computed the uncertainties by calculating the sample variance of the different forward passes using Equation 3. We built single-task models using the saved model weights to predict the target labels: income for ADULT, IHM for MIMIC-III, and the 4 mood-related labels for SNAPSHOT. With the predictions obtained from these models, we calculated the DIRs. These steps provided us with distributions of 100 performance scores and 100 DIR scores. We implemented the non-dominated sorting algorithm introduced in 3.4 to obtain the sets of Pareto optimal solutions for the experiments for each dataset. This algorithm computed an output containing the models with the best trade-off between the 2 scores (performance and DIR), and for which there were no better-performing models. Any model can be chosen from these models according to what is being prioritized for the particular problem type and dataset. We plotted the performance scores against the DIR scores to get a visual representation of their relationship.
We also conducted further analysis to determine whether or not there is a relationship between uncertainty and fairness. For this, we explored the relationship between the uncertainty scores and fairness scores (DI ratio) of the model ensemble produced by the Monte Carlo method.
5.3 Reweighing Method Implementation
We compared our proposed method to a standard bias mitigation method called Reweighing. Reweighing is a pre-processing bias mitigation technique. It involves applying appropriate weights to different tuples in the training dataset to make it discrimination free with respect to the protected label Park et al. (2021). It generates weights for the training examples in each (group, label) combination differently to ensure fairness before classification. Bellamy et al. (2018). See the details of Reweighting in Appendix B.
We applied the AI Fairness 360 implementation of this method to our processed data for all datasets and passed the weight vectors to our baseline model for all 3 datasets and calculated their fairness metric scores to compare to our proposed method.

### v2

@@ 3 Fairness Terminology
In this section, we introduce and define some of the concepts and terminologies generally used in algorithmic fairness research that are relevant to this paper.
@@
  • Disparate impact ratio (DIR): This is a fairness metric which compares the rate of positive outcomes in the unprivileged class, with the rate of positive outcomes in the privileged class, as shown in Equation 1. It is the measure of how different outcomes are for different groups, based on the results of a model (Feldman et al. (2015)). A value of less than 1 indicates less favorable outcomes for the unprivileged group. In this paper, we assume an acceptable lower bound of 0.8, and an upper bound of 1.2, with 1 being the ideal score.
@@
  • Equalized odds: This fairness metric enforces that the model correctly identifies the positive outcome at equal rates across groups, and miss-classify the positive outcome at equal rates across groups (creating the same proportion of True Positives and False Positives across groups). In this work, we represent this metric using 2 values; Difference in False Negative (FN) scores, and Difference in False Positive (FP) scores.
4 Multi-task Learning Based Bias Mitigation Method (Proposed Method)
4.1 Overview
Our proposed methodology is based on the premise of Bayesian uncertainty estimation to address fairness in machine learning models. Uncertainty estimation refers to estimating the confidence level of a model output. In this case, the posterior variance represents the total uncertainty of the prediction (Wang et al. (2021b)). In our proposed method, we hypothesize that when the model exhibits the greatest range of uncertainty regarding the protected label, it indicates a deficiency in knowledge within specific areas of the feature space linked to that label. Utilizing the model weights from this range of uncertainty for predicting the target label is likely to diminish the influence of the protected label on the final prediction outcome. This approach is based on our findings that a spectrum of model weights, rather than a single optimal model, yields the best fairness scores. Therefore, we employ non-dominated sorting to identify the Pareto optimal set of solutions. This method will allow us to select models that effectively balance fairness and performance by recognizing that improvements in one metric may lead to reductions in the other, thus providing a robust framework for addressing fairness. Figure 1) represents a visual diagram of our proposed method, the following subsections detail the flow of our method with the different concepts involved.
@@ 4.1 Overview
4.2 Multi-Task Learning Framework
To accurately assess the uncertainty associated with both the target and protected labels, we first implement an MTL framework that optimizes the learning of correlated tasks by sharing common features while allowing for task-specific adaptations. This framework uses a shared input layer but distinct output layers for each task—one for the primary target prediction and another for protected label prediction. This framework is rooted in the principle that related tasks can share useful information and thus be learned more effectively together than separately (Caruana (1997)).
4.3 Monte Carlo Dropout for Uncertainty Estimation
Our MTL model integrates Monte Carlo Dropout for uncertainty estimation. Unlike traditional dropout techniques used solely during training to combat overfitting, MC Dropout extends this concept to inference. We incorporate MC Dropout in our MTL model by performing multiple forward passes with randomly applied dropout. This simulates drawing samples from the model’s posterior weight distribution. This process effectively turns our neural network into a Bayesian model, where each dropout configuration leads to different sets of predictions. From this, we get predictions for both the target and the selected protected label. Analyzing the variance across these predictions provides a robust measure of uncertainty, particularly highlighting where the model’s predictions are least confident, potentially signaling underlying biases.
Modeling uncertainty with MC Dropout works by running multiple forward passes through the model with a different dropout mask every time. It draws weights from approximate posterior distributions. The posterior distribution is a way to summarize what we know about uncertain quantities in Bayesian analysis. It combines the prior distribution and the likelihood function, which tells you what information is in your observed data (the “new evidence”). In other words, the posterior distribution summarizes what we know after the data is observed.
@@ 4.3 Monte Carlo Dropout for Uncertainty Estimation
Further details on the theoretical underpinnings of MC Dropout can be found in the seminal work by Gal and Ghahramani (2016). The uncertainties derived from the Monte Carlo simulations are critical for assessing model fairness.
@@ 4.4 Non-Dominated Sorting
From the saved models obtained using Monte-Carlo Dropout, we compute the uncertainty scores and extract the parameters from each model. These parameters are then employed in a distinct (non-multi-task) model dedicated to predicting our main target label. We train these models, monitor their performance metrics, and compute fairness metrics, including DIRs, differences in FN and FP rates from their predictions.
To identify the most effective models in terms of balancing performance and fairness, we apply a non-dominated sorting algorithm to the ensemble’s performance and fairness scores, particularly focusing on the DIRs. This process allows us to derive the Pareto optimal set of models, collectively referred to as the "Pareto Front." Models on this front represent the optimal trade-offs between performance and fairness; any improvements in one metric would necessitate a compromise in the other. This front is significant as it contains no models that outperform others across both metrics simultaneously, illustrating the inherent trade-offs between achieving high performance and maintaining fairness.
The non-dominated sorting algorithm operates by comparing models based on their performance scores (X) and fairness scores (Y - specifically DIR scores), as outlined in Algorithm 1. It ranks models that lie on the Pareto frontier, sorted in order of their optimality. The algorithm’s objective is to maximize the performance scores (X) while minimizing the absolute difference between the fairness scores (Y) and the ideal value of 1. It identifies Pareto optimal points from two metric arrays $Xs$ and $Ys$, which represent the conflicting objectives of performance and fairness. The algorithm pairs these metrics, sorts them based on the $Xs$ values according to the preference for maximization or minimization, and iteratively constructs a Pareto frontier. Each subsequent pair of X and Y is evaluated against the last point on the frontier: it is added if it offers an improvement in $Y$ without worsening $X$, measured by the absolute difference from an ideal value. The outcome is two arrays, $p\_frontX$ and $p\_frontY$, depicting the optimal trade-offs between the evaluated metrics.
Algorithm 1 Calculate Pareto Frontier
1:procedure ParetoFrontier($Xs,Ys,maxX,maxY$)
2:     Input: Metric arrays: Performance scores ($Xs$), Fairness scores ($Ys$) of equal length; booleans $maxX$, $maxY$
3:     Output: Arrays $p\_frontX$, $p\_frontY$ of Pareto optimal points
4:
5:     Combine $Xs$ and $Ys$ into list $myList$ of tuples $(x,y)$
6:     if $maxX$ is True then
7:         Sort $myList$ by $x$ in descending order
8:     else
9:         Sort $myList$ by $x$ in ascending order
10:     end if
11:     Initialize $p\_front$ with the first tuple from $myList$
12:     for each tuple $(currentX,currentY)$ in $myList$ do
13:         if $maxY$ then
14:              if $\text{abs}(1-currentY)<\text{abs}(1-y$ of last tuple in $p\_front)$ then
15:                  Append $(currentX,currentY)$ to $p\_front$
16:              end if
17:         else
18:              if $\text{abs}(1-currentY)>\text{abs}(1-y$ of last tuple in $p\_front)$ then
19:                  Append $(currentX,currentY)$ to $p\_front$
20:              end if
21:         end if
22:     end for
23:     Separate $p\_front$ into $p\_frontX$ and $p\_frontY$
24:     return $(p\_frontX,p\_frontY)$
25:end procedure
5 Experiments
5.1 Datasets
@@
We evaluate the effectiveness of our proposed method using three datasets, (1) ADULT (individual income and demographic information) (Kohavi et al. (1996)), (2) MIMIC-III (data from patients in critical care units) Johnson et al. (2016), and (3) SNAPSHOT datasets (multi-modal physiological, behavioral, and survey data collected from 350 college students) (Sano et al. (2015)). Table 1 describes the input and binary target variables for each of these datasets used in our experiments. More information on the details of these datasets and data processing can be found in Appendix A.
Table 1: Input and Target Variables for ADULT, MIMIC-III, and SNAPSHOT Datasets
Dataset Input Variables Target Variables
ADULT Age, workclass, education, Salary
Marital-status, occupation,
relationship, race, sex
MIMIC-III Vital signs, medications, In-hospital mortality (IHM)
laboratory measurements,
fluid balance, procedure codes,
hospital length of stay,
survival data,
laboratory test results,
diagnostic codes,
patient demographics,
billing information,
observations and notes, etc.
SNAPSHOT Physiological features, Morning happiness
mobile phone usage, Evening happiness
weather data Morning Calmness
Evening Calmness
For the ADULT dataset, we regard age, sex(gender) and race as the protected labels, and for MIMIC-III, age, gender, insurance type, and marital status. For the SNAPSHOT dataset, we assign gender, race, and ethnicity as the protected labels in our experiments. We also use personality types, including openness, conscientiousness, extraversion, agreeableness, and neuroticism, as protected labels in our experiments for SNAPSHOT. Table 2 details how we binary-encode each protected label and assign them as privileged and unprivileged classes to analyze them using our chosen fairness metrics.
@@
Table 2: Binarization of Protected labels for ADULT, MIMIC-III, and SNAPSHOT Datasets
Dataset Label Privileged Class (1) Unprivileged Class (0)
ADULT Age Less than or equal to 40 years Greater than 40 years
@@
MIMIC-III Age Greater than 60 years Less than or equal to 60 years
@@
SNAP-SHOT Gender Male Female
Race White Non-white
Ethnicity Non-Hispanic/latino Hispanic/latino
Openness $<=50$ $>50$
Conscientiousness $>50$ $<=50$
Extraversion $<=50$ $>50$
Agreeableness $<=50$ $>50$
Neuroticism $<=50$ $>50$
@@ 5.2 Baseline Models and Fairness Analyses
The first part of our experiments involves developing baseline models that make the target predictions for all datasets (income for ADULT, IHM for MIMIC-III, and 4 happiness and calmness labels for SNAPSHOT). Next, we analyze our prediction results for fairness by calculating the DIR and equalized odds (FP and FN rates) for predictions against the protected labels for all datasets. We conduct these analyses to determine whether and where the models were introducing bias and ensure that we only apply our bias mitigation method to aspects of the data that are actually biased. Details of the models can be found in Appendix B.
After performing fairness analysis on each baseline model, we select the protected labels by which our models showed the most bias and test our methods on all datasets. For all fairness analyses in our experiments, we utilize a fairness toolkit developed by IBM - AI Fairness 360 (Bellamy et al. (2019)).
@@ 5.3 Proposed Method Implementation
We apply our bias-reducing MTL method by developing MTL models to predict target and protected labels by which the baseline models are biased. We first train the MTL model for each of the combinations for all datasets. We assign different loss weight ratios to the label pairs, prioritizing the prediction label over the protected label to ensure that the model prioritizes the prediction label and makes certain that its uncertainty score is kept lower than the protected label prediction. We train the MTL networks for 100 epochs for all the datasets and save the weights at each epoch, making it a total of 100 weight combinations for each experiment run.
After running the MTL networks and obtaining an ensemble of models, we calculate the uncertainties of these models. We computed the uncertainties by calculating the sample variance of the different forward passes using Equation 3. We build single-task models using the saved model weights to predict the target labels: income for ADULT, IHM for MIMIC-III, and the 4 mood-related labels for SNAPSHOT. With the predictions obtained from these models, we calculate the DIRs. These steps provide us with distributions of 100 performance scores and 100 DIR scores. We implement the non-dominated sorting algorithm introduced in 4.4 to obtain the sets of Pareto optimal solutions for the experiments for each dataset. This algorithm computes an output containing the models with the best trade-off between the performance score and the fairness score (DIR), and for which there were no better-performing models. Any model can be chosen from these models according to what is being prioritized for the particular problem type and dataset. We plot the performance scores against the DIR scores to get a visual representation of their relationship.
5.4 Benchmark Bias Mitigation Approaches
We compare our proposed method to the following standard and state of the art bias mitigation methods applied in prior work.
(1) Reweighting (Kamiran and Calders (2012)) (ADULT, MIMIC-III, and SNAPSHOT): a standard bias mitigation method that has been used extensively in the field and is considered a benchmark technique for bias mitigation.
(2) Adversarial Reweighted Learning (ARL) (Lahoti et al. (2020)) (ADULT and MIMIC-III): uses an adversarial setup to reweight features for fairness, which presents a direct method to manipulate feature importance based on their impact on bias, while our approach, using MC Dropout, offers a probabilistic view of feature importance by assessing uncertainty.
(3) FairRF (Zhao et al. (2022)) (ADULT and MIMIC-III): focuses on mitigating bias through feature manipulation, and provides a direct contrast to our method’s use of uncertainty estimation to influence model predictions. This comparison aims to illustrate how managing uncertainty can potentially be more effective than directly modifying feature weights.
Comparing these approaches can highlight the benefits of a probabilistic versus adversarial framework in bias mitigation. See details of each baseline model in Appendix B.